# ELICIT: LLM Augmentation via External In-Context Capability

### v1

@@ Abstract
Enhancing the adaptive capabilities of large language models is a critical pursuit in both research and application. Traditional fine-tuning methods require substantial data and computational resources, especially for enhancing specific capabilities, while in-context learning is limited by the need for appropriate demonstrations and efficient token usage. Inspired by the expression of in-context learned capabilities through task vectors and the concept of modularization, we propose ELICIT, a framework consisting of two modules designed to effectively store and reuse task vectors to elicit the diverse capabilities of models without additional training or inference tokens. Our comprehensive experiments and analysis demonstrate that our pipeline is highly transferable across different input formats, tasks, and model architectures. ELICIT serves as a plug-and-play performance booster to enable adaptive elicitation of model capabilities. By externally storing and reusing vectors that represent in-context learned capabilities, ELICIT not only demonstrates the potential to operate modular capabilities but also significantly enhances the performance, versatility, adaptability, and scalability of large language models. Our code will be publicly available  [^1] https://github.com/LINs-lab/ELICIT.
@@ 1 Introduction
  • Selective Capability Activation: ELICIT allows for targeted activation of specific model capabilities based on the input query In our experiments with a math-only capability library, ELICIT boosted Math performance dramatically (e.g., from $2.6$% to $21.3$% for Mistral) while maintaining or slightly improving performance in other domains (Section 4.3).
  • Complementary Integration: ELICIT shows potential for complementary use with existing methods, offering further performance gains. When combined with BM25 retrieval, ELICIT enhanced Pythia’s average performance from $21.9$% to $30.1$% (Section 4.5).
@@ 2 Related Work
Inspired by findings that intermediate representations in LLMs encode semantic meaning (Zou et al., 2023), researchers have explored injecting in-context learning demonstrations, encoded as function vectors, into intermediate representations to trigger desired predictions (Liu et al., 2023b; Hendel et al., 2023; Todd et al., 2023; Li et al., 2024). However, this line of work focuses on manipulating internal representations. We are the first to comprehensively explore the modular approach of externally storing and retrieving such task representations to augment large language model capabilities.
@@ 3.2 Building Capabilities Library
  • We obtain the task vector $\boldsymbol{\theta}\in\mathbb{R}^{L\times d}$ by processing the ICL prompt $\mathbf{p}_{i}^{t}$ (defined in (1), using $N=16$ demonstrations). This is done by extracting the hidden states $\{\mathbf{h}_{l}\}_{l=1}^{L}$ from all layers of the language model. We also store these ICL prompts for future use.
@@ 4.2 Efficient Capability Elicitation
Table 2: Performance of ELICIT across model and tasks. ELICIT significantly enhances performance while maintaining the same token usage as Zero-shot, often achieving results comparable to or better than 16-shot and BM25 ICL retriever methods. This improvement is consistent across various models and tasks, demonstrating ELICIT’s efficiency and effectiveness in boosting model capabilities without increasing computational demands.
@@
16-shot 1574.7 ± 0.2 42.4 ± 0.9 36.1 ± 0.4 59.9 ± 0.2 21.1 ± 1.9 50.9 ± 0.6 42.1 ± 0.2
bm25 2145.5 ± 29.1 44.3 ± 1.4 47.1 ± 0.3 61.6 ± 0.8 25.7 ± 1.8 45.6 ± 2.1 44.9 ± 0.4
Zero-shot 108.2 ± 1.4 32.2 ± 1.2 32.9 ± 0.2 42.5 ± 1.2 14.0 ± 1.0 35.5 ± 1.2 31.4 ± 0.7
Llama3 ELICIT 108.2 ± 1.4 38.1 ± 0.9 46.1 ± 0.3 60.7 ± 1.2 19.4 ± 1.1 49.4 ± 2.1 42.7 ± 0.8
16-shot 1797.3 ± 2.8 50.0 ± 0.9 41.6 ± 0.5 60.8 ± 0.9 17.6 ± 1.6 55.5 ± 1.3 45.1 ± 0.5
bm25 2403.5 ± 32.1 44.5 ± 1.3 43.1 ± 0.4 60.1 ± 1.0 18.6 ± 0.8 52.3 ± 2.6 43.7 ± 0.7
Zero-shot 123.4 ± 1.7 29.5 ± 1.2 26.2 ± 0.4 45.5 ± 1.3 2.8 ± 0.1 36.1 ± 0.3 28.0 ± 0.5
Mistral ELICIT 123.4 ± 1.7 36.9 ± 1.0 47.1 ± 0.6 60.4 ± 1.1 20.4 ± 1.1 47.9 ± 1.6 42.5 ± 0.5
16-shot 1597.6 ± 2.4 48.1 ± 0.4 22.2 ± 0.8 12.5 ± 0.7 10.3 ± 0.9 28.2 ± 0.9 24.3 ± 0.4
bm25 2190.5 ± 30.2 38.8 ± 0.6 20.3 ± 0.1 12.5 ± 0.6 11.7 ± 0.8 26.2 ± 1.3 21.9 ± 0.2
Zero-shot 109.8 ± 1.5 42.9 ± 0.3 18.9 ± 0.2 21.8 ± 1.6 7.3 ± 0.2 32.5 ± 1.2 24.7 ± 0.4
Pythia ELICIT 109.8 ± 1.5 60.1 ± 0.1 25.7 ± 0.9 20.9 ± 1.2 14.4 ± 1.3 40.9 ± 2.5 32.4 ± 0.4
16-shot 1601.5 ± 1.7 31.0 ± 0.9 24.0 ± 0.7 27.2 ± 1.1 12.6 ± 1.0 36.4 ± 0.4 26.2 ± 0.3
bm25 2190.5 ± 30.2 34.6 ± 1.8 23.3 ± 0.6 23.8 ± 1.1 12.9 ± 2.1 29.3 ± 1.6 24.8 ± 0.3
Zero-shot 109.8 ± 1.5 36.1 ± 1.5 19.5 ± 0.5 17.3 ± 1.2 5.8 ± 1.2 30.2 ± 0.2 21.8 ± 0.2
Mamba ELICIT 109.8 ± 1.5 45.6 ± 1.4 29.3 ± 0.7 30.9 ± 1.4 15.4 ± 1.4 47.3 ± 1.6 33.7 ± 0.7
@@
From Tables 2, comparing the zero-shot baseline and ELICIT, we observe that ELICIT significantly elicits model capabilities across most tasks without increasing token usage. Across the 20 tasks, ELICIT achieves an average improvement of 11.4% across different models. For Llama3, ELICIT improves over zero-shot by 15.1% while using the same 108.2 tokens. ELICIT demonstrates substantial gains in Reasoning (e.g., +14.4% for LLama3) and Safety tasks (e.g., +14.0% for LLama3). In many cases, ELICIT’s performance is comparable to or surpasses that of 16-shot and BM25 methods, despite their higher token requirements. Furthermore, it exhibits robustness across various template formats, highlighting its versatility.
@@ 4.3 Selective Adaptive Activation of Capabilities
We demonstrate selective activation by constructing a library containing only math-related task vectors, as shown in Figure 5. The results clearly illustrate that ELICIT significantly boosts performance in the Math domain, with a dramatic increase from $2.6$% to $21.3$%, while maintaining performance in other domains for Mistral. Notably, the Reasoning domain also shows a slight improvement, increasing from $22.3$% to $22.4$%. Results for other models, presented in Appendix F, demonstrate a similar trend. The striking improvement in Math performance, coupled with the subtle gain in Reasoning and the stability in other domains, demonstrates ELICIT’s capacity for targeted capability activation, making it a flexible and efficient performance enhancer.
@@ 4.4 ELICIT generalize to Unseen Tasks without addtional information
In Table 3, we observe that ELICIT significantly improves model performance on unseen tasks (GLUE-COLA, BBQ Religion, Deepmind (Saxton et al., 2019), MMLU-Psychology, and BBH-Logical-Deduction-Five-objects) not present in its capability library. Across all models, ELICIT consistently outperforms the Zero-shot baseline. In several cases, it even approaches or surpasses the BM25 retrieval baseline, despite using substantially fewer tokens. For instance, on the MMLU-Psychology task, ELICIT achieves a $21.7$% absolute improvement over Zero-shot, surpassing BM25 retrieval by $5.7$% for Llama3 Model. These results are achieved without additional token usage and task information, demonstrating ELICIT’s efficiency, flexibility, and generalization ability.
Table 3: ELICIT can generalize to unseen tasks. ELICIT achieves significant performance gains without additional token usage across different models and unseen tasks.
@@
BM25 2066.9 ± 22.5 83.6 ± 0.8 54.5 ± 0.7 29.1 ± 1.9 81.6 ± 0.4 29.6 ± 0.0 55.7 ± 0.2
Zero-shot 102.9 ± 0.7 72.0 ± 0.7 38.6 ± 1.1 17.5 ± 2.6 54.2 ± 0.3 14.3 ± 0.0 39.3 ± 0.8
Llama ELICIT 102.9 ± 0.7 73.4 ± 0.6 47.8 ± 0.8 23.1 ± 3.4 75.9 ± 0.2 26.2 ± 0.0 49.3 ± 0.5
BM25 2300.3 ± 23.5 84.6 ± 1.2 57.6 ± 0.6 19.5 ± 2.2 74.7 ± 0.9 15.1 ± 0.0 50.3 ± 0.5
Zero-shot 114.6 ± 0.8 43.3 ± 1.1 35.4 ± 3.3 9.0 ± 0.4 57.9 ± 0.7 6.9 ± 0.0 30.5 ± 1.0
Mistral ELICIT 114.6 ± 0.8 46.2 ± 2.0 45.0 ± 2.7 25.4 ± 1.7 71.2 ± 0.4 24.0 ± 0.0 42.4 ± 1.1
BM25 2140.4 ± 23.8 67.4 ± 0.7 30.4 ± 2.9 19.6 ± 1.3 15.9 ± 0.4 14.6 ± 0.0 29.6 ± 0.8
Zero-shot 105.9 ± 0.7 70.3 ± 1.0 21.3 ± 2.9 10.9 ± 0.7 13.9 ± 0.5 6.2 ± 0.0 24.5 ± 0.4
Pythia ELICIT 105.9 ± 0.7 72.4 ± 0.4 33.2 ± 2.7 23.3 ± 1.7 26.4 ± 0.4 13.9 ± 0.0 33.8 ± 0.4
BM25 2140.4 ± 23.8 54.3 ± 1.3 29.7 ± 0.5 13.1 ± 2.0 15.1 ± 0.3 5.8 ± 0.0 23.6 ± 0.4
Zero-shot 105.9 ± 0.7 48.5 ± 0.6 22.0 ± 2.0 9.8 ± 1.3 19.9 ± 0.9 10.0 ± 0.0 22.0 ± 0.2
Mamba ELICIT 105.9 ± 0.7 51.5 ± 1.0 31.0 ± 4.5 13.5 ± 3.0 24.1 ± 1.6 18.3 ± 0.0 27.7 ± 1.1
@@ 4.5 Complementary Intergration
ELICIT serves as a versatile plug-and-play performance booster.
Tables 4 and 9 showcase ELICIT’s remarkable ability to function as an orthogonal, plug-and-play enhancement to existing solutions like BM25 retrieval. In Table 4, we observe consistent performance gains across all models and tasks when ELICIT is combined with BM25 retrieval. Notably, Pythia demonstrates substantial improvements, with its average performance increasing from 21.9% to 30.1% for in-domain tasks, and from 23.6% to 28.4% on unseen tasks. For instance, the Llama3 model on the MMLU-Psychology task achieves $81.2$% with ELICIT, surpassing BM25 retreival alone ($81.6$%), persisting across various models and tasks. The combination with BM24 Retrieval results on unseen tasks is shown in H. Its plug-and-play nature allows for seamless integration with existing methods, enhancing model performance in both in-domain and unseen task scenarios without compromising the benefits of current approaches.
Table 4: ELICIT as an plug-and-play performance booster: Performance Gains When Combined with BM25 on In-Domain Tasks.
@@
Llama BM25 44.3 ± 1.3 47.1 ± 0.3 61.6 ± 0.8 25.7 ± 1.8 45.6 ± 2.1 44.9 ± 0.4
BM25+ELICIT 43.9 ± 1.7 49.9 ± 0.5 60.5 ± 1.0 25.2 ± 1.2 48.7 ± 3.2 45.6 ± 0.7
Mistral BM25 44.5 ± 1.3 43.1 ± 0.4 60.0 ± 1.0 18.6 ± 0.8 52.3 ± 2.6 43.7 ± 0.8
BM25+ELICIT 47.7 ± 1.3 47.1 ± 0.6 61.8 ± 1.5 23.5 ± 2.2 49.2 ± 3.0 45.9 ± 0.2
Pythia BM25 38.8 ± 0.7 20.4 ± 0.2 12.6 ± 0.6 11.7 ± 0.7 26.2 ± 1.3 21.9 ± 0.2
BM25+ELICIT 60.9 ± 1.0 23.4 ± 1.0 15.4 ± 0.8 14.3 ± 1.5 36.3 ± 0.9 30.1 ± 0.3
Mamba BM25 34.6 ± 1.8 23.4 ± 0.6 23.7 ± 1.1 12.9 ± 2.0 29.3 ± 1.6 24.8 ± 0.3
BM25+ELICIT 44.4 ± 1.9 31.1 ± 0.3 30.4 ± 1.7 15.9 ± 2.7 42.3 ± 2.1 32.8 ± 0.3

### v2

@@ Abstract
Enhancing the adaptive capabilities of large language models is a critical pursuit in both research and application. Traditional fine-tuning methods require substantial data and computational resources, especially for enhancing specific capabilities, while in-context learning is limited by the need for appropriate demonstrations and efficient token usage. Inspired by the expression of in-context learned capabilities through task vectors and the concept of modularization, we propose ELICIT, a framework consisting of two modules designed to effectively store and reuse task vectors to elicit the diverse capabilities of models without additional training or inference tokens. Our comprehensive experiments and analysis demonstrate that our pipeline is highly transferable across different input formats, tasks, and model architectures. ELICIT serves as a plug-and-play performance booster to enable adaptive elicitation of model capabilities. By externally storing and reusing vectors that represent in-context learned capabilities, ELICIT not only demonstrates the potential to operate modular capabilities but also significantly enhances the performance, versatility, adaptability, and scalability of large language models. Our code is publicly available  [^1] https://github.com/LINs-lab/ELICIT.
@@ 1 Introduction
  • Selective Capability Activation: ELICIT allows for targeted activation of specific model capabilities based on the input query. In our experiments with a math-only capability library, ELICIT boosted Math performance dramatically (e.g., from $2.6$% to $21.3$% for Mistral) while maintaining or slightly improving performance in other domains (Section 4.3).
  • Complementary Integration: ELICIT shows potential for complementary use with existing methods, offering further performance gains. When combined with BM25 retrieval, ELICIT enhanced Pythia’s average performance from $22.1$% to $28.3$% (Section 4.5).
@@ 2 Related Work
Inspired by findings that intermediate representations in LLMs encode semantic meaning (Zou et al., 2023), researchers have explored injecting in-context learning demonstrations, encoded as function vectors, into intermediate representations to trigger desired predictions (Liu et al., 2023b; Hendel et al., 2023; Todd et al., 2023; Li et al., 2024). The scope of this research line has broadened to include different modalities,with recent work demonstrating its effectiveness in both visual (Hojel et al., 2025) and multimodal domains (Huang et al., 2024). However, this line of work focuses on manipulating internal representations. We are the first to comprehensively explore the modular approach of externally storing and retrieving such task representations to augment large language model capabilities.
@@ 3.2 Building Capabilities Library
  • We obtain the task vector $\boldsymbol{\theta}\in\mathbb{R}^{L\times d}$ by processing the ICL prompt $\mathbf{p}_{i}^{t}$ (defined in (1), using randomly selected $N=16$ demonstrations). This is done by extracting the hidden states $\{\mathbf{h}_{l}\}_{l=1}^{L}$ from all layers of the language model. We also store these ICL prompts for future use.
@@ 4.2 Efficient Capability Elicitation
Table 2: Performance of ELICIT across model and tasks. ELICIT significantly enhances performance while maintaining the same token usage as Zero-shot, often achieving results comparable to or better than 16-shot and 16-shot BM25 ICL retriever methods. This improvement is consistent across various models and tasks, demonstrating ELICIT’s efficiency and effectiveness in boosting model capabilities without increasing computational demands. We sample 100 examples per task across three random seeds.
@@
16-shot 1883.8 ± 0.9 60.6 ± 1.0 56.0 ± 0.4 70.6 ± 1.0 26.7 ± 2.0 62.1 ± 0.4 55.2 ± 0.4
bm25 2350.7 ± 24.9 56.1 ± 1.5 68.8 ± 0.2 69.5 ± 0.9 28.0 ± 2.3 56.7 ± 2.0 55.8 ± 0.7
Zero-shot 108.3 ± 1.4 32.2 ± 1.2 31.6 ± 0.2 42.5 ± 1.2 14.0 ± 1.0 35.5 ± 1.2 31.2 ± 0.7
Llama3 ELICIT 108.3 ± 1.4 41.6 ± 0.4 46.7 ± 0.1 60.6 ± 1.4 19.1 ± 1.4 49.9 ± 2.1 43.5 ± 0.8
16-shot 2161.3 ± 0.9 55.3 ± 0.5 52.1 ± 0.5 70.8 ± 0.4 23.7 ± 1.7 63.1 ± 0.6 53.0 ± 0.1
bm25 2655.2 ± 27.3 55.2 ± 0.3 66.0 ± 0.5 70.2 ± 1.9 24.1 ± 0.4 62.1 ± 0.5 55.5 ± 0.4
Zero-shot 123.5 ± 1.7 29.6 ± 1.2 26.9 ± 0.4 45.5 ± 1.3 2.8 ± 0.1 36.1 ± 0.3 28.2 ± 0.5
Mistral ELICIT 123.5 ± 1.7 41.9 ± 1.0 48.3 ± 0.3 59.4 ± 0.9 20.3 ± 0.9 48.7 ± 1.8 43.7 ± 0.6
16-shot 1942.4 ± 0.9 50.2 ± 0.5 19.6 ± 0.1 12.8 ± 0.9 9.2 ± 1.6 31.8 ± 0.9 24.7 ± 0.2
bm25 2422.8 ± 26.0 33.3 ± 2.2 25.8 ± 0.4 12.9 ± 0.5 11.0 ± 1.8 27.3 ± 2.1 22.1 ± 0.5
Zero-shot 110.0 ± 1.5 43.0 ± 0.4 18.3 ± 0.3 22.0 ± 1.5 7.3 ± 0.1 32.5 ± 1.2 24.6 ± 0.4
Pythia ELICIT 110.0 ± 1.5 64.0 ± 1.6 23.6 ± 1.1 20.4 ± 1.4 14.5 ± 1.0 41.2 ± 2.5 32.7 ± 0.5
16-shot 1942.4 ± 0.9 37.5 ± 1.0 31.5 ± 0.5 31.6 ± 0.8 14.2 ± 0.5 41.7 ± 1.2 31.3 ± 0.3
bm25 2422.8 ± 26.0 29.3 ± 2.2 34.9 ± 0.9 24.7 ± 0.5 15.1 ± 2.2 35.4 ± 1.2 27.9 ± 0.3
Zero-shot 110.0 ± 1.5 36.1 ± 1.5 19.3 ± 0.5 17.3 ± 1.2 5.8 ± 1.2 30.1 ± 0.1 21.7 ± 0.2
Mamba ELICIT 110.0 ± 1.5 51.1 ± 0.7 28.7 ± 0.8 29.2 ± 1.3 15.3 ± 1.1 48.2 ± 1.8 34.5 ± 0.6
@@
From Tables 2, comparing the zero-shot baseline and ELICIT, we observe that ELICIT significantly elicits model capabilities across most tasks without increasing token usage. Across the 20 tasks, ELICIT achieves an average improvement of 11.4% across different models. For Llama3, ELICIT improves over zero-shot by 12.3% while using the same 108.2 tokens. ELICIT demonstrates substantial gains in Reasoning (e.g., +15.1% for LLama3) and Safety tasks (e.g., +14.4% for LLama3). In some cases, ELICIT’s performance is comparable to or surpasses that of 16-shot and BM25 methods, despite their higher token requirements. Furthermore, it exhibits robustness across various template formats, highlighting its versatility.
@@ 4.3 Selective Adaptive Activation of Capabilities
We demonstrate selective activation by constructing a library containing only math-related task vectors, as shown in Figure 5. The results clearly illustrate that ELICIT significantly boosts performance in the Math domain, with a dramatic increase from $2.6$% to $21.3$%, while maintaining performance in other domains for Mistral. Notably, the Reasoning domain also shows a slight improvement, increasing from $22.3$% to $22.4$%. This behavior stems from ELICIT’s selective application of task vectors from library, which are not applied when no relevant tasks vectors are detected. More discussion is presented in Appendix J. Results for other models, presented in Appendix F, demonstrate a similar trend. The striking improvement in Math performance, coupled with the subtle gain in Reasoning and the stability in other domains, demonstrates ELICIT’s capacity for targeted capability activation, making it a flexible and efficient performance enhancer.
@@ 4.4 ELICIT generalize to Unseen Tasks without addtional information
In Table 3, we observe that ELICIT significantly improves model performance on unseen tasks (GLUE-COLA, BBQ Religion, Deepmind (Saxton et al., 2019), MMLU-Psychology, and BBH-Logical-Deduction-Five-objects) not present in its capability library. Across all models, ELICIT consistently outperforms the Zero-shot baseline. In several cases, it even approaches or surpasses the BM25 retrieval baseline, despite using substantially fewer tokens. For instance, on the MMLU-Psychology task, ELICIT achieves a $15.8$% absolute improvement over Zero-shot for Llama3 Model. These results are achieved without additional token usage and task information, demonstrating ELICIT’s efficiency, flexibility, and generalization ability.
Table 3: ELICIT can generalize to unseen tasks. ELICIT achieves significant performance gains without additional token usage across different models and unseen tasks. We sample 100 examples per task across three random seeds. We use BM25 retrieval of 16 examples as baseline.
@@
BM25 2502.8 ± 26.0 55.4 ± 1.0 64.6 ± 1.3 30.7 ± 1.7 83.0 ± 0.1 48.3 ± 0.0 56.4 ± 0.4
Zero-shot 103.6 ± 47.7 72.0 ± 0.7 38.6 ± 1.1 17.5 ± 2.6 54.2 ± 0.3 17.1 ± 0.0 39.9 ± 0.8
Llama ELICIT 103.6 ± 47.7 63.4 ± 0.9 45.0 ± 0.7 23.7 ± 3.4 70.0 ± 0.6 25.7 ± 0.0 45.6 ± 0.4
BM25 2804.6 ± 27.6 44.4 ± 2.2 70.7 ± 0.7 26.6 ± 3.9 78.7 ± 1.1 25.7 ± 0.0 49.2 ± 0.3
Zero-shot 115.4 ± 51.0 43.3 ± 1.1 35.4 ± 3.3 9.0 ± 0.4 57.9 ± 0.7 7.4 ± 0.0 30.6 ± 1.0
Mistral ELICIT 115.4 ± 51.0 41.7 ± 0.8 42.1 ± 2.5 25.1 ± 1.2 65.6 ± 0.6 15.6 ± 0.0 38.0 ± 0.6
BM25 2600.0 ± 28.3 5.8 ± 1.0 19.1 ± 1.2 14.1 ± 1.2 4.7 ± 0.3 1.0 ± 0.0 8.9 ± 0.3
Zero-shot 106.7 ± 49.6 48.5 ± 0.6 21.7 ± 1.7 9.7 ± 1.2 20.1 ± 0.8 7.6 ± 0.0 21.5 ± 0.1
Pythia ELICIT 106.7 ± 49.6 45.4 ± 0.6 30.3 ± 4.2 14.2 ± 1.8 20.4 ± 0.6 14.3 ± 0.0 24.9 ± 0.6
BM25 2600.0 ± 28.3 48.1 ± 3.1 30.6 ± 1.1 21.6 ± 3.3 19.1 ± 0.9 25.8 ± 0.0 29.0 ± 0.9
Zero-shot 106.7 ± 49.6 70.3 ± 1.0 21.3 ± 2.9 10.9 ± 0.7 13.9 ± 0.5 6.2 ± 0.0 24.5 ± 0.4
Mamba ELICIT 106.7 ± 49.6 63.6 ± 0.4 31.5 ± 2.5 22.1 ± 3.3 20.4 ± 0.2 14.4 ± 0.0 30.4 ± 0.9
@@ 4.5 Complementary Intergration
ELICIT shows potential as a plug-and-play performance booster.
While ELICIT demonstrates compatibility with existing solutions like BM25 retrieval, Table 4 reveals nuanced performance patterns. For smaller models (Pythia-2.8B and Mamba-2.8B), combining ELICIT with BM25 yields consistent improvements, with Pythia’s average performance increasing from 22.1% to 28.3% (+5.9%). However, larger models (Llama3-8B and Mistral-7B) exhibit mixed results: while NLU and Reasoning tasks show modest gains (e.g., +2.6% for Llama3), Knowledge and Safety tasks experience slight declines. Aligning with the findings of Li et al. (2024), this phenomenon can be attributed to two factors: (1) smaller models’ relatively weak in-context learning capabilities benefit more from additional task-relevant information provided by our method, while (2) larger models’ inherently stronger in-context adaptation abilities may be disrupted by the introduction of additional context that alters their learned representations. Future work could investigate this scale-dependent phenomenon.
Table 4: ELICIT as a potential plug-and-play performance booster: performance when combined with BM25 on in-domain tasks. Results indicate stronger complementary effects for smaller models (Pythia, Mamba), while larger models (Llama3, Mistral) show task-specific variations.
@@
Llama BM25 56.1 ± 1.5 68.8 ± 0.2 69.5 ± 0.9 28.0 ± 2.3 56.7 ± 2.0 55.8 ± 0.7
BM25+ELICIT 58.0 ± 0.4 62.7 ± 0.5 65.1 ± 0.5 25.1 ± 1.6 54.5 ± 3.5 53.1 ± 1.1
Mistral BM25 55.2 ± 0.3 66.0 ± 0.5 70.2 ± 1.9 24.1 ± 0.4 62.1 ± 0.5 55.5 ± 0.4
BM25+ELICIT 54.5 ± 0.8 62.6 ± 0.4 67.5 ± 1.7 24.8 ± 1.9 58.0 ± 1.4 53.5 ± 0.5
Pythia BM25 33.3 ± 2.2 25.8 ± 0.4 12.9 ± 0.5 11.0 ± 1.8 27.3 ± 2.1 22.1 ± 0.5
BM25+ELICIT 53.5 ± 1.5 26.5 ± 1.1 14.5 ± 0.6 13.2 ± 1.7 33.7 ± 0.7 28.3 ± 0.3
Mamba BM25 29.3 ± 2.2 34.9 ± 0.9 24.7 ± 0.5 15.1 ± 2.2 35.4 ± 1.2 27.9 ± 0.3
BM25+ELICIT 38.4 ± 1.2 31.6 ± 0.4 28.9 ± 0.2 15.2 ± 3.8 42.9 ± 1.9 31.4 ± 0.3